OpenStudy (anonymous):

4b-6n=-36 3b-6n=-36 elimination using multiplacation. solve for x and y.

OpenStudy (kuoministers):

you know how to use the elimination method? subtract or add vertically?

OpenStudy (anonymous):

multiplication for this one.

OpenStudy (kuoministers):

i dont think you can multiply in elimination?

OpenStudy (kuoministers):

in this you can add vertically

OpenStudy (anonymous):

its the only way to get the right answer with this problem. its math hw.

OpenStudy (kuoministers):

4b-6n=-36 3b-6n=-36 7b =0

OpenStudy (kuoministers):

so you can figure it out now?

OpenStudy (anonymous):

7b=0 doesnt make sense. o/7=0.

OpenStudy (kuoministers):

b = 0 n =6

OpenStudy (anonymous):

ok....i'll be checking later. thanks.